OpenStudy (donnie1999):

Marge runs an ice cream parlor. Her specialty is triple chocolate sundaes. She can prepare 1 sundae every 2 minutes, and she earns $1.20 for each sundae she makes. If she just makes sundaes for a single shift of at most 4 hours and at least 2 hours, which function relates her earnings to the number of minutes she works?

OpenStudy (donnie1999):

\[a. f(x)=1.02 \left| x/2 \right|, if 2 \le x \le 4\]

OpenStudy (donnie1999):

\[b. f(x)=1.2\left| x/2 \right|,if 120\le x \le 240\]

OpenStudy (donnie1999):

\[c. f(x)=2.4 \left| x \right|.if 120 \le x \le 240\]

OpenStudy (donnie1999):

\[d. f(x) = 1/2 \left| 1.2 \right|, if 120\le x \le 240\]

OpenStudy (donnie1999):

\[d. f(x) = 1/2 \left| 1.2 \right|, if 120\le x \le 240\]

OpenStudy (donnie1999):

\[d. f(x) = 1/2 \left| 1.2 \right|, if 120\le x \le 240\]

OpenStudy (butterflydreamer):

remember, you want to relate earnings to number of minutes she works. So we know she works for 2 - 4 hrs right? So we need to convert 2 and 4 hrs to minutes. Also we know she earns $1.20 for each sundae she makes. She takes 2 minutes to make 1 sundae... therefore if x = sundae, 1.20 ( x/2) ^ here we have earning times how many sundaes she can produce per minute.

OpenStudy (donnie1999):

@butterflydreamer i got the answer b. but i am not sure if that's wright.

OpenStudy (butterflydreamer):

that should be correct :)